# Revealing the Geometrical and Vibrational Properties of the Defects Driving the Boson Peak

### v1

@@ Abstract
The vibrational density of states is key to understanding the mechanical, thermal, and transport properties of materials. In amorphous solids, this density shows an excess of vibrational modes compared to the Debye model, known as the boson peak, whose origin remains poorly understood. Previous studies have suggested a link between the boson peak and quasi-localized nonphononic vibrations, or "defects." However, it has been difficult to clearly identify these defects, possibly because they hybridize with extended phonons, casting doubt on their existence and connection to the boson peak. In this work, we introduce a simple and practical method for separating hybridized phonons from localized vibrations. We show that phonons at the boson peak frequency hybridize with localized defects. These defects are anisotropic, compact, and exhibit oscillatory pure shear deformations. Their density correlates with the excess of vibrational modes at the boson peak frequency across various two- and three-dimensional systems, confirming that they are the microscopic origin of the boson peak.
@@
Thermal and vibrational properties of amorphous solids display universal characteristics that are very different from their crystalline counterparts [1]. In particular, their vibrational density of states (vDOS) $D(\omega)$ presents a universal excess anomaly over the Debye prediction $D(\omega)\propto\omega^{D-1}$, with $D$ the number of spatial dimensions and $\omega$ the frequency, known as the “boson peak” (BP). The microscopic origin of the BP remains one of the most controversial topics in condensed matter physics.
@@
Some theories assume that BP arises from additional nonphononic excitations associated with quasi-localised modes (QLMs). The observation that structural glasses present nonphononic excitations is old [2]. This idea was inspired (see e.g. [3]) by the concept of resonant modes first formulated in the context of defective crystals [4] and it is a recurrent theme within the soft potential model of glassy anomalies [5, 6, 7, 8]. These nonphonic excitations have been widely reported in the low-frequency spectrum of amorphous solids and even disordered crystals [9], where they contribute to the density of states with a faster than Debye term, $D(\omega)\propto\omega^{p}$ with $p>(D-1)$, [10, 11, 12]. They have a quadrupolar nature and are characterized by a core made of few atoms accompanied by long-range displacements with power-law decay [13]. It has been argued that QLMs also exist at the boson peak frequency and cause the boson peak  [14, 7, 8, 15, 16, 17]. However, this scenario is difficult to prove, as QLMs would strongly hybridize with extended phononic modes around the BP frequency, rendering their identification [6, 18] and the study of their properties problematic. Henceforth, the existence of localised modes at the boson peak is questionable, motivating the development of alternative theories of the BP [19, 20, 21, 22, 23, 24, 25].
@@
Recent works have explored the possibility of disentangling QLMs and extended modes at the boson peak frequency. In one approach, deformations with seemingly quadrupolar structures were studied within the vibrational modes at the boson peak frequency to assess whether they could serve as seeds for energy-minimal localized deformations [16, 17], which are identified with QLMs. Alternatively, Refs.[25, 26] examined how modes near the boson peak contribute to the amplitude of particle vibrational motion (see also[3]). Particles receiving significant contributions from these modes tend to form clusters, possibly emerging as a consequence of the hybridization of phonons and QLMs, and referred to as stringlets [25, 26] due to their resemblance to one-dimensional string-like or ring-like dynamical defects. Furthermore, the interaction between stringlets and phonons induces a flat boson peak mode [27, 28], observed in both simulations [25] and experiments [29]. This supports the hypothesis that stringlets, rather than quasi-localized quadrupolar modes, may be the microscopic origin of the boson peak, even in heated crystals [30]. It remains unclear whether localized defects are responsible for the boson peak in glasses and whether these defects are quasi-localized quadrupolar modes or string-like excitations, leaving the microscopic origin of the boson peak unresolved.
@@
In this work, we resolve this tension by introducing a simple and efficient method based on cage-relative displacement to disentangle extended phonons from quasi-localized defects. We demonstrate that the modes at the boson peak frequency are a superposition of phonons and anisotropic, compact localized defects, exhibiting pure shear oscillatory deformations in two and three spatial dimensions. We show that the density of these defects explains the excess number of modes at the boson peak, providing strong support for previous suggestions that the boson peak originates from quadrupolar quasi-localized excitations [15, 16]. Additionally, we identify the so-called stringlets [25], revealing that, contrary to earlier speculations, they are not string-like objects but compact ones that align with the localized defects we have uncovered. Our work unifies different theoretical perspectives on the boson peak in glasses, resolving the long-standing debate and potentially establishing a consensus on its microscopic origin.
Figure 1: Attraction regulates the boson peak in 2D and 3D amorphous solids. a The reduced vDOS in three and (inset) two dimensions for systems differing in the extension $x_{c}$ of the intermolecular attractive potential well. In 3D, the materials lose stability as $x_{c}$ increases, as the boson peak intensity increases (b) while its frequency decreases (c). In 2D, attraction does not sensibly influence the excess vibrational modes. Frequencies are expressed in units of the natural frequency $\omega_{0}=c_{s}/a_{0}$, with $a_{0}$ the interparticle spacing and $c_{s}$ the shear wave speed. The inset of (c) illustrates the localised defects that, according to our investigation, induce the boson peak in a three-dimensional system with $x_{c}=1.7$.
Models and vibrational density of states
@@
We numerically investigate the vibrational properties of two- and three-dimensional systems of polydisperse particles interacting via a Lennard-Jones-like potential $U(r_{ij},x_{c})$ which vanishes with its first two derivatives at $x_{c}\sigma$, $\sigma$ being the average particle diameter [31, 32]. We provide details in Methods. Previous works have shown that $x_{c}$, which controls the extension of the attractive well, influences the system’s relaxation dynamics [32] and the mechanical response [31, 33, 34, 35]. We simulate $N=256k$ particle systems in square or cubic simulation boxes with periodic boundary conditions at fixed interparticle spacing $a_{0}=\rho^{-1/d}\simeq 0.977$. We generate amorphous solid configurations by minimizing, via conjugate gradient, the energy of systems in thermal equilibrium in the NVT ensemble at $T=4.0\epsilon$, above the glass transition temperature for the considered $x_{c}$ values [32]. All data presented below are averaged over $200$ configurations at each $x_{c}$.
We evaluate the vibrational density of states (vDOS) over a broad frequency range by performing a Fourier transform of the velocity autocorrelation function in the linear response regime. Additionally, we compute the vDOS by directly diagonalizing the Hessian matrix, allowing us to access the eigenvectors, ${\bf e}_{k}$, and eigenfrequencies, $\omega_{k}$, in a low-frequency regime that includes the boson peak (BP) frequency. The two methods give consistent results (see SI). Fig. 1(a) illustrates the reduced-vDOS $D(\omega)/A_{D}\omega^{D}$, with $D_{\rm Debye}=A_{D}\omega^{D}$ Debye’s prediction (see SI), and demonstrates the BP in both two and three spatial dimensions. In 3D, the normalized peak amplitude increases with $x_{c}$ while its frequency $\omega_{\rm bp}$ decreases, indicating that shorter attractive wells lead to more stable glasses, as previously suggested [15, 33, 35, 31]. Due to the increase in stability limit we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$, phonons band structure persists to high frequency given our considered system sizes, making $\omega_{\rm bp}$ of difficult identification. In 2D, the role of $x_{c}$ is minor. In this case, we consider $x_{c}\geq 1.4\sigma$; at smaller $x_{c}$ the BP peak becomes very broad and $\omega_{\rm bp}$ of difficult estimation.
Disentangling phonons and localised defects
Due to the interaction with phonons, localised vibrations cannot be directly visualized as harmonic vibrations when their frequency is inside a phonon band. Instead, the phonons-defect hybridization leads to extended harmonic modes combining the two. Devising strategies to disentangle this hybridization and expose the localised defects is a longstanding challenge. Two approaches have been introduced so far: one relies on the assumption that the localised defects appear in regions where the local vibrational kinetic energy is high [3, 25, 26], and the other on the hypothesis that these defects have a quadrupolar structure visible in the hybridized field [16, 17].
@@
We propose an alternative, simple and practical approach to disentangle localised defects and extended plane waves based on the hypothesis that plane waves at the BP frequency are locally affine while localised defects are not. Diverse approaches could be employed to identify regions of local non-affine behaviour [36]. Here we rely on the cage-relative displacement field, which has been previously used to filter out the effect of long-wavelength plane waves [37, 38, 39], to separate and remove the plane-wave contribution from the eigenvectors. This approach involves analyzing particles’ displacements relative to their immediate surroundings or the ‘cage’ formed by neighbouring particles. Specifically, given a displacement field ${\bm{u}}$, the cage-relative displacement ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}$ of particle $i$ is ${\bm{u}}^{\leavevmode\nobreak\ \rm cr}_{i}={\bm{u}}_{i}-\frac{1}{N_{v}}\sum_{j=1}^{N_{v}}{\bm{u}}_{j}$, where the summation extends over the $N_{v}$ Voronoi neighbors of particle $i$. We use this approach by considering as displacement fields the vibrational modes of the systems, ${\bm{e}}_{k}$, which we determine via the direct diagonalization of the Hessian matrix. We consider as defective particles those with an $|{\bm{e}}^{\leavevmode\nobreak\ \rm cr}_{k,i}|^{2}$ value in the top 5%, unless otherwise specified, and determine the vibrational defects via a cluster analysis of these particles, considering two particles as belonging to the same cluster if Voronoi neighbors.
@@
We validate our approach in Fig. 2. First, we demonstrate that this method does not filter out localized vibrations. In panel (a), we present a snapshot of a two-dimensional system where each particle is colour-coded based on its $|{\bm{e}}_{k,i}|^{2}$, for a low-frequency localised mode $k$. The localised excitation remains clearly visible in panel (b), where particles are coloured according to the magnitude of their cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$, proving that our approach preserves localized defects. In panels (c) and (d), we compare the magnitude of the displacement and cage-relative displacement for the mode at the boson peak frequency. The cage-relative measure effectively filters out the collective translations responsible for the extended clusters in panel (c), revealing defective regions of large non-affine particle motion. This approach works equally well in three-dimensional systems, as shown in the inset of Fig. 1(c). Additional evidence in the Supplementary Information (SI) shows that cage-relative measures successfully filter out plane wave contributions at the boson peak frequency.
@@
Figure 2: Disentangling extended plane waves and localised defects. a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with a low-frequency mode with $\omega_{k}=0.23\omega_{\rm bp}$. A localized defect is visible. b The defect remains visible in the colour map of the magnitude of the cage-relative (CR) displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$. Panels c and d use the same approach to visualize the mode at the BP frequency. In this case, the cage-relative measure reveals the presence of many vibrational defects not apparent in the standard measure. In all plots, the magnitudes are scaled so that their average value is $1$.
Geometrical properties of vibrational defects
Fig. 3(a) shows that the probability distribution of the number of particles in a cluster $P(n)$ decays abruptly at small $n$ and roughly exponentially at larger $n$, similar to the distribution of the number of particles participating in the stringlets [40, 41, 25]. On increasing $x_{c}$, the average cluster size varies from $\langle n\rangle=16$ to $26$, in $3D$, and $\langle n\rangle=9$ to $13$, in $2D$. The average coordination number $z(n)$ grows with $n$ due to the competition between bulk and surface particles, as in Fig. 3(b), and it is larger than two if not for minute clusters, indicating that the clusters are not string-like. To assess a cluster’s geometrical properties, we perform a principal component analysis of the covariance matrix $C_{\alpha\beta}=\langle r_{\alpha}r_{\beta}\rangle-\langle r_{\alpha}\rangle\langle r_{\beta}\rangle$ of the positions of its particles. We use the eigenvalues $\lambda_{i}$ of $C_{\alpha\beta}$, which measure the variances of the particle positions along the principal directions, to define normalized standard deviations $\sigma_{i}=\sqrt{\lambda_{i}}/\sum_{i}\sqrt{\lambda_{i}}$ that obey the geometrical constraint $\sum_{i}\sigma_{i}=1$, and order them such that $\sigma_{i+1}>\sigma_{i}$. In two dimensions, $\sigma_{1}=1$ for linear (maximally anisotropic, one-dimensional) clusters and $\sigma_{1}=\frac{1}{2}$ for circular (isotropic) clusters. We find that the distribution $P(\sigma_{1})$ peaks at $\sigma_{1}\simeq 0.65$ and vanishes for $\sigma_{1}>0.8$, as shown in Fig. 3(c). This result further confirms that defects are not one-dimensional but rather asymmetric two-dimensional structures.
In three dimensions, for linear (1D), flat (2D), and spherical (isotropic 3D) clusters, $\sigma_{1}=1$, $\sigma_{1}+\sigma_{2}=1$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=\frac{1}{3}$, respectively. Fig. 3(d) presents density maps of the $P(\sigma_{1},\sigma_{2})$ probability distribution at $x_{c}=1.4$ and $x_{c}=1.7$ (shifted vertically) and suggests the localised defects have three-dimensional geometry with a variable degree of anisotropy.
@@
Figure 3: Geometrical properties of vibrational defects. a Probability distribution of the number of particles belonging to each defect in 3D and 2D (inset). Symbols identify the values of $x_{c}$, as in Fig. 1b. b Average coordination number of the particles in clusters of size $n$. c We normalise ($\sum\sigma_{i}=1$) and order ($\sigma_{i}>\sigma_{i+1}$) the standard deviations of the particles’ positions along the principal axis of each ‘defect’, considering defects with at least $n=3$ particles. In two dimensions, the probability distribution of $P(\sigma_{1})$ peaks at intermediate values and drops at large ones, indicating that the defects are not one-dimensional ($\sigma_{1}=1$), but two-dimensional objects, albeit not precisely circular $(\sigma_{1}=1/2)$. d In three dimensions, the probability distribution $P(\sigma_{1},\sigma_{2})$ clarifies that clusters are not one-dimensional ($\sigma_{1}=1$) or two-dimensional $(\sigma_{1}+\sigma_{2}=1)$. We thus assume they are three-dimensional, albeit non-spherical $(\sigma_{1}=\sigma_{2}=1/3)$. Data for $x_{c}=1.7$ are shifted vertically for clarity.
Quadrupolar nature of vibrational defects
We obtain an insight into the vibrational nature of these defects by considering that the cage-relative displacement is conceptually related to a local strain tensor, as they are both insensitive to local uniform translations. We thus evaluate the local strain tensor ${\bm{\epsilon}}=\frac{1}{2}\left({\bm{\nabla}}{\bm{e}}+({\bm{\nabla}}\bm{e})^{T}\right),$ associated with the displacement field (eigenvector, ${\bm{e}}$) of the mode at the boson peak frequency, coarse-grained on grids of side-length $w\simeq 1.1$. We decompose this tensor into its volumetric and deviatoric components: $\bm{\varepsilon}=\bm{\varepsilon}_{\rm vol}+\bm{\varepsilon}_{\rm dev}$. $\bm{\varepsilon}_{\rm vol}$ represents local volumetric deformations. In contrast, the traceless deviatoric component represents volume-preserving deformations, which dominate the vibrational spectrum at the boson peak [42]. The eigenvalues $\upsilon_{k}$ of $\bm{\varepsilon}_{\rm dev}$ satisfy the traceless condition $\sum\upsilon_{k}=0$ and define a local quadrupolar charge $Q_{\rm loc}=\left(\frac{1}{2}\sum_{k=1}^{D}\upsilon_{k}^{2}\right)^{1/2}$ [43, 44].
Fig. 4(a) demonstrates a high Pearson correlation coefficient between the local quadrupolar charge and the coarse-grained cage-relative displacement magnitude, consistently with the expected analogy between local strain and cage-relative measure. Since defects are clusters of particles with significant coarse-grained cage-relative displacement, we calculate the deviatoric strain of a defect by summing the deviatoric strain of the individual particles within it. Specifically, we define a cluster deviatoric strain as $\mathcal{E}_{\rm dev}=\sum\bm{\varepsilon}_{\rm dev}$, where the sum extends over all bins containing particles of the considered cluster. The eigenvectors ${\bf U}_{k}$ and eigenvalues $\Upsilon_{k}$ of $\mathcal{E}_{\rm dev}$ describe the vibrational motion of the cluster. A cluster extends into the directions with $\Upsilon_{k}>0$ while it shrinks in the others for half period of oscillation, with particles reversing their direction of motion during the remaining half period. In the following, we assume $\Upsilon_{1}<0$ and order the eigenvalues so that $|\Upsilon_{i}|>|\Upsilon_{i+1}|$.
In two dimensions, the traceless condition implies $\Upsilon_{1}=-\Upsilon_{2}$. Figure 4(b) illustrates the quadrupoles associated with each cluster on a heat map of the magnitude of the coarse-grained cage-relative displacement, finding expected correlations in light of panel (a). The figure illustrates a small fraction of the system and represents each quadrupole with four arrows with a length proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$.
In three dimensions, we examine the relevance of various deformation modes that satisfy the traceless condition $\sum\Upsilon_{i}=0$ by analyzing the probability distribution $P(\Upsilon_{1},\Upsilon_{2})$. This distribution is supported within a range defined by the line $\Upsilon_{1}+\Upsilon_{2}=0$, which represents pure shear deformation, and the line $\Upsilon_{1}+2\Upsilon_{2}=0$, which corresponds to volume-preserving isotropic ($\Upsilon_{3}=\Upsilon_{2}=-\Upsilon_{1}/2$) uniaxial compressions. Figure 4(c) shows that quadrupolar deformations are predominantly pure shear, as $P(\Upsilon_{1},\Upsilon_{2})$ peaks near the $\Upsilon_{1}+\Upsilon_{2}=0$ line for different values of $x_{c}$. This result is further supported by direct visualization of the quadrupoles (Figure 4(d)), depicted as collections of 6 arrows with lengths proportional to ±$\pm\Upsilon_{i}{\bm{U}}_{i}$. Since $|\Upsilon_{3}|\ll|\Upsilon_{2}|$, the blue arrows corresponding to displacements in the ${\bm{U}_{3}}$ direction are not visible, making the quadrupoles appear as a two-dimensional cross.

### v2

# Geometrical and Vibrational Properties of the Defects Driving the Boson Peak

@@ Abstract
In amorphous solids, the vibrational density of states shows an excess of modes over the Debye model, known as the boson peak, whose origin remains unclear. Studies suggest a link to quasi-localized nonphononic vibrations or ‘defects,’ but identifying them is challenging due to hybridization with phonons that renders methods based on localization properties, such as the participation ratio, unreliable. We introduce a practical method to separate hybridized phonons from localized vibrations and find that boson peak phonons hybridize with compact, two-dimensional defects exhibiting oscillatory pure shear deformations. These two-dimensional defects are also exposed by the procedure recently employed to identify stringlets (Nature Physics volume 18, pages 669–677 (2022)), suggesting that these may not be one-dimensional objects as speculated. Our work demonstrates the presence of localized defects at the boson peak frequency and provides a comprehensive characterization of their vibrational and geometric properties, resolving the tension between the concepts of quasi-localized quadrupolar defects and stringlets.
@@
Thermal and vibrational properties of amorphous solids display universal characteristics that are very different from their crystalline counterparts [1]. In particular, their vibrational density of states (vDOS) $D(\omega)$ presents a universal excess anomaly over the Debye prediction $\color[rgb]{0,0,0}D(\omega)\propto\omega^{d-1}$}, with $\color[rgb]{0,0,0}d$} the number of spatial dimensions and $\omega$ the frequency, known as the “boson peak” (BP). The microscopic origin of the BP remains one of the most controversial topics in condensed matter physics.
@@
Some theories assume that BP arises from additional nonphononic excitations associated with quasi-localised modes (QLMs). The observation that structural glasses present nonphononic excitations is old [2]. This idea was inspired (see e.g. [3, 4]) by the concept of resonant modes first formulated in the context of defective crystals [5] and it is a recurrent theme within the soft potential model of glassy anomalies [6, 7, 8, 9]. These nonphononic excitations have been widely reported in the low-frequency spectrum of amorphous solids and even disordered crystals [10], where they contribute to the density of states with a faster than Debye term, $D(\omega)\propto\omega^{p}$ with $\color[rgb]{0,0,0}p>(d-1)$}[11, 12, 13, 14, 15]. They have a quadrupolar nature and are characterized by a core made of few atoms accompanied by long-range displacements with power-law decay [16]. It has been argued that similar defects also exist at higher frequencies and cause the boson peak [17, 8, 9, 18, 19, 20]. However, this scenario is difficult to prove, as these defects would strongly hybridize with extended phononic modes around the BP frequency. In particular, methods based on localization properties, such as the participation ratio, are not effective at these frequencies. If defect exists at high frequencies, they do not appear as QLM, rendering their identification [7, 21] and the study of their properties problematic. Henceforth, the existence of localised defects at the boson peak is questionable, motivating the development of alternative theories of the BP [22, 23, 24, 25, 26, 27, 28].
@@
Recent works have explored the possibility of disentangling localised defects and extended modes at the boson peak frequency. In one approach, deformations with seemingly quadrupolar structures were visually isolated from vibrational modes at the boson peak frequency to assess whether they could serve as seeds for energy-minimal localized deformations [19, 20], which are identified with QLMs. However, local quadrupolar deformations are not generally apparent in the hybridized displacement field, as demonstrated in Appendix A, and defects might, in principle, induce diverse localized deformations. Alternatively, one could examine how modes near the boson peak contribute to the amplitude of particle vibrational motion [3]. Refs. [28, 29] have suggested that particles receiving significant contributions from these modes tend to organize along string-like clusters or stringlets with one-dimensional nature and lacking quadrupolar characteristics. The interaction between stringlets and phonons induces a flat boson peak mode [30, 31], observed in both simulations [28] and experiments [32] supports the hypothesis that stringlets, rather than quasi-localized quadrupolar modes, could be the microscopic origin of the boson peak, even in heated crystals [33]. It remains unclear whether localized defects are responsible for the boson peak in glasses and whether these defects are quasi-localized quadrupolar modes or string-like excitations. Thus, the microscopic origin of the boson peak remains unresolved.
@@
In this work, we resolve this tension by introducing a simple and efficient method based on cage-relative (CR) displacement to disentangle extended phonons from quasi-localized defects. We demonstrate that the modes at the boson peak frequency are a superposition of phonons and anisotropic, compact, two-dimensional localized defects, exhibiting pure shear oscillatory deformations in two and three spatial dimensions. We show that the size of these defects correlates with the length scales associated with the boson peak, providing strong support for previous suggestions that the boson peak originates from quadrupolar quasi-localized excitations [18, 19]. Additionally, we identify the so-called stringlets [28], revealing that, contrary to earlier speculations, they are not string-like objects but compact ones that align with the localized defects we have uncovered. Our work unifies different theoretical perspectives on the boson peak in glasses, resolving the long-standing debate and potentially establishing a consensus on its microscopic origin.
@@
We numerically investigate the vibrational properties of two- and three-dimensional systems of polydisperse particles interacting via a Lennard-Jones-like potential $U(r_{ij},x_{c})$ which vanishes with its first two derivatives at $x_{c}\sigma$, $\sigma$ being the average particle diameter of the interacting particles [34, 35]. Previous works have shown that $x_{c}$, which controls the extension of the attractive well, influences the system’s relaxation dynamics [35] and the mechanical response [34, 36, 37, 38]. We simulate systems with $N=32k$ particles - which are enough for finite size systems to be negligible at the boson peak frequency [39, 29, 19] - in square or cubic simulation boxes with periodic boundary conditions at fixed interparticle spacing $a_{0}=\rho^{-1/D}\simeq 0.977$. We generate amorphous solid configurations by minimizing, via conjugate gradient, the energy of systems in thermal equilibrium in the NVT ensemble at $T=4.0\epsilon$, above the glass transition temperature for the considered $x_{c}$ values [35]. All data presented below are averaged over $200$ configurations at each $x_{c}$. We provide details on the model, the numerical procedures, and the $x_{c}$ dependence of the vibrational model in Appendix B.
@@
Exposing hybridized defects – We propose a simple and practical approach to disentangle localised defects and extended plane waves based on the hypothesis that plane waves at the BP frequency are locally affine while localised defects are not. Diverse approaches could be employed to decompose a vibration into a phononic and a non-phononic component [40, 41], or more generally, to identify regions of local non-affine behaviour [42, 43]. Here, we rely on the cage-relative (CR) displacement field to separate and remove the plane-wave contribution from the eigenvectors. This approach has been previously used to filter out the effect of low-frequency long-wavelength plane waves on the low-temperature liquid dynamics [44, 45, 46], never to isolate defects. This approach involves analyzing particles’ displacements relative to their immediate surroundings or the ‘cage’ formed by neighbouring particles. Specifically, given a displacement field $\bm{u}$}, the cage-relative displacement ${\bm{u}}^{\rm cr}_{i}$ of particle $i$ is ${\bm{u}}^{\rm cr}_{i}={\bm{u}}_{i}-\frac{1}{N_{v}}\sum_{j=1}^{N_{v}}{\bm{u}}_{j}$, where the summation extends over the $N_{v}$ voronoi neighbors of particle $i$. We use this approach by considering as displacement fields the vibrational modes of the systems, ${\bm{e}}_{k}$, which we determine through the direct diagonalization of the Hessian matrix.
@@
We validate this approach in Appendix C by demonstrating that the CR measure preserves the localized defects and successfully filters out the plane wave contributions at the boson peak frequency. As an example, Fig. 1 illustrates a snapshot of a two-dimensional system where each particle is colour-coded based on the magnitude of the displacement $|{\bm{e}}_{k,i}|^{2}$ (in panel a), and cage-relative displacement $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$ (in panel b), for an eigenmode at the boson peak frequency. The cage-relative measure reveals defective regions of large non-affine particle motion. This approach works equally well in three-dimensional systems, as shown in the inset of Fig. 6(c).
We consider as defective particles those with an $|{\bm{e}}^{\penalty 10000\ \rm cr}_{k,i}|^{2}$ value in the top 5%, unless otherwise specified, and determine the vibrational defects via a cluster analysis of these particles, considering two particles as belonging to the same cluster if voronoi neighbours. In the following, we discuss results obtained via the analysis of a single mode at the BP frequency, for each system. We have checked that analogous results occur for close frequencies $|\omega-\omega_{\rm bp}|<0.1\omega_{\rm bp}$, but not at higher ones.
@@
Figure 1: a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with an eigenmode with $\omega_{k}=\omega_{\rm bp}$. Swirls characterizing hybridization with plane waves are observed. b The cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$ reveals the presence of many vibrational defects not apparent in the standard measure. In these plots, the magnitudes are scaled so that their average value is 11.
@@
Figure 2: a Probability distribution of the number of particles belonging to each defect in 3D and 2D (inset). b Average coordination number of the particles in clusters of size $n$. c We normalise ($\sum\sigma_{i}=1$) and order ($\sigma_{i}>\sigma_{i+1}$) the standard deviations of the particles’ positions along the principal axis of each ‘defect’, considering defects with at least $n=3$ particles. In two dimensions, the probability distribution of $P(\sigma_{1})$ peaks at intermediate values and drops at large ones, indicating that the defects are not one-dimensional ($\sigma_{1}=1$), but two-dimensional objects, albeit not precisely circular $(\sigma_{1}=1/2)$. d In three dimensions, the probability distribution $P(\sigma_{1},\sigma_{2})$ clarifies that clusters are not one-dimensional ($\sigma_{1}=1$) or two-dimensional $(\sigma_{1}+\sigma_{2}=1)$. We thus assume they are three-dimensional, albeit non-spherical $(\sigma_{1}=\sigma_{2}=1/3)$. Data for $x_{c}=1.7$ are shifted vertically for clarity. e Radius of gyration, $R_{g}$, as a function of cluster size $n$. 3D data are shifted by a factor 33 for clarity. f The characteristic length scale $\lambda$ of defects governs the boson peak frequency in two and three dimensions. In all panels, symbols identify the values of $x_{c}$.
Geometrical properties – Fig. 2(a) shows that the probability distribution of the number of particles in a cluster $P(n)$ decays abruptly at small $n$ and roughly exponentially at larger $n$, similar to the distribution of the number of particles participating in the stringlets [47, 48, 28]. On increasing $x_{c}$, the average cluster size varies from $\langle n\rangle=16$ to $26$, in 3D, and $\langle n\rangle=9$ to $13$, in 2D. The average coordination number $\color[rgb]{0,0,0}Z(n)$} grows with $n$ due to the competition between bulk and surface particles, as in Fig. 2(b), and it is larger than two if not for minute clusters, indicating that the clusters are not string-like.
To further assess a cluster’s geometrical properties, we perform a principal component analysis of the covariance matrix $C_{\alpha\beta}=\langle r_{\alpha}r_{\beta}\rangle-\langle r_{\alpha}\rangle\langle r_{\beta}\rangle$ of the positions of its particles. We use the eigenvalues $\lambda_{i}$ of $C_{\alpha\beta}$, which measure the variances of the particle positions along the principal directions, to define normalized standard deviations $\sigma_{i}=\sqrt{\lambda_{i}}/\sum_{i}\sqrt{\lambda_{i}}$ that obey the geometrical constraint $\sum_{i}\sigma_{i}=1$, and order them such that $\sigma_{i}>\sigma_{i+1}$. In two dimensions, $\sigma_{1}=1$ for linear (maximally anisotropic, one-dimensional) clusters and $\sigma_{1}=\frac{1}{2}$ for circular (isotropic) clusters. We find that the distribution $P(\sigma_{1})$ peaks at $\sigma_{1}\simeq 0.65$ and vanishes for $\sigma_{1}>0.8$, as shown in Fig. 2(c). In three dimensions, for linear (1D), flat (2D), and spherical (isotropic 3D) clusters, $\sigma_{1}=1$, $\sigma_{1}+\sigma_{2}=1$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=\frac{1}{3}$, respectively. Fig. 2(d) presents density maps of the $P(\sigma_{1},\sigma_{2})$ probability distribution at $x_{c}=1.5$ and $x_{c}=1.7$ (shifted vertically) and suggests that the localised defects are three-dimensional compact (panel b) clusters with a high degree of anisotropy. Due to this anisotropy, the radius of gyration scales as $R_{g}\propto n^{q}$, with $q\simeq 0.55$ in 2D and $q\simeq 0.5$ in 3D, if not for large clusters, as shown in Fig. 2(e). The distribution of the clusters’ radius of gyration is exponentially distributed (see Appendix D), allowing us to associate a typical length scale $\lambda$ to each $x_{c}$. Fig. 2(f) shows that $\omega_{\rm bp}\propto c_{s}/\lambda$, in both two and three dimensions, proving these defects’ relevance to the boson peak phenomenology.